OpenStudy (wade123):

@Jhannybean @ganeshie8 can you guys please explain this calc problem?!

OpenStudy (wade123):

just number 5!!

Nnesha (nnesha):

calculator problem ???

OpenStudy (anonymous):

calculus darling :P

Nnesha (nnesha):

ohh okay ik

OpenStudy (anonymous):

sure u do ;)

ganeshie8 (ganeshie8):

where are you stuck ? just integrate the absolute value of velocity for the distance

ganeshie8 (ganeshie8):

\[\int\limits_0^8 |v(t)|~dt\]

OpenStudy (calculusfunctions):

First find the position function s(t)by integrating the velocity function as @ganeshie8 has suggested. Then the average velocity on the given interval will be found by\[\frac{ s(8)-s(0) }{ 8-0}\]

OpenStudy (wade123):

integrating: t^3/3-9t^t/2+18t+C , what do i do next?

ganeshie8 (ganeshie8):

#1 should make you think about average value of a funciton thats precicely the reason we define average value of f(x) in interval [a, b] as \[\large \dfrac{\int\limits_a^b f(x)~dx}{b-a}\]

OpenStudy (wade123):

yrah i did those, i only havnt done 5

OpenStudy (wade123):

just 5!

ganeshie8 (ganeshie8):

for #5, start by finding the absolute value of v(t)

ganeshie8 (ganeshie8):

set v(t) equal to 0 and solve t

OpenStudy (wade123):

x=6 and x=3

ganeshie8 (ganeshie8):

Excellent! next split the integral and work each of them

OpenStudy (wade123):

like plug them in?

OpenStudy (wade123):

i dont get what that means:/

OpenStudy (calculusfunctions):

What he means his now find the distance from t = 0 to t = 3, then from t = 3 to t = 6, then from t = 6 to t = 8. Then the sum of all these distances will be the total distance travelled from t = 0 to t = 8.

ganeshie8 (ganeshie8):

Notice v(t) is positive everywhere except in the interval [3, 6] your goal is to integrate |v(t)|, so you may split the integral into 3 pieces such that the v(t) doesn't change sign in each interval: \[\int\limits_0^8 |v(t)|~dt = \left|\int\limits_0^3 v(t) ~dt \right|+ \left|\int\limits_3^6 v(t) ~dt \right|+ \left|\int\limits_6^8 v(t) ~dt \right| \]

OpenStudy (wade123):

and solve each one?

ganeshie8 (ganeshie8):

yes, you may use the antiderivative that you have found earlier and simply work the bounds if you want

OpenStudy (wade123):

thank you(:

ganeshie8 (ganeshie8):

you're welcome!

OpenStudy (calculusfunctions):

If you already have the position function, which is\[s(t) = \frac{ 1 }{ 3 }t ^{3}-\frac{ 9 }{ 2 }t ^{2}+18t +1\]Then for example the distance from t = 0 to t = 3 is\[|s(3)-s(0)|\]